# Representation Learning of Multivariate Time Series using Attention and Adversarial Training

### v1

@@ 3.0.4 CAE:
    This model shows low similarity (DTW) and low diversity (Entropy) compared to the other models. The low diversity is further illustrated by the distribution in the t-SNE representation in Figure 2. Similar to TAE-GAN, the distribution lies exclusively within a new mode; there is no overlap with the modes of the dataset as the latent space was not regularized any further. The generated time series in Figure 7 in the appendix in Section 0.A.2 confirm the low similarity to the dataset and the low diversity. In contrast to the observations with TAE, the test error is higher than CAE with GAN training.
@@ 3.0.5 CAE-GAN:
    The Entropy score shows a high diversity compared to the other models, also confirmed by the distribution in the t-SNE plot (Figure 2). Here, the artificial time series lie within the different modes of the dataset, although not sufficiently distributed. The similarity score (DTW) is still low in an overall comparison, however slightly increased compared to CAE. The error plots in Figure 12 in the appendix in Section 0.A.2 show that the regulation by the GAN training was very unstable. There was no steady learning effect of the discriminator; only in isolated epochs it learned to clearly distinguish between true and fake samples, while for the most part it was completely uncertain (outputs near to $0.5$). This possibly resulted in hardly useful gradients for the generator. The generated time series in Figure 13 in the appendix in Section 0.A.2 affirms the low similarity to the dataset.
@@ 3.0.6 CAE-WGAN:
    With an Entropy of $0.855$ this model shows the highest diversity, which is also confirmed by the distribution in the t-SNE plot (Figure 2). The generated time series are well balanced among the different modes of the dataset. Nevertheless, the similarity (DTW) remained low and even worsened compared to CAE and CAE-GAN. The GAN training has also been very unstable (Figure 14 in the appendix in Section 0.A.2); However, a continuous learning effect occurs in the last 250 epochs. Therefore, a longer training might have led to convergence. Like CAE-GAN, the generated time series in Figure 15 in the appendix in Section 0.A.2 do not show similar patterns to the dataset.
@@ 4 Conclusion
    In this work, a Transformer-based model is proposed to generate multivariate time series signals within the support of the underlying distribution of an existing dataset. The Transformer architecture was augmented with a bottleneck to act as an autoencoder and was optimized using a conventional GAN and the Wasserstein GAN approach. Additionally, the model was compared to an existing approach that uses a convolutional autoencoder for local neighborhood generation of time series signals. The results show that the Transformer model was more suitable for stable GAN training. Using the Wasserstein GAN approach, the model outperformed the other models in terms of similarity to the dataset, however not all modes were sufficiently represented. The diversity is estimated to be rather low based on the multivariate Entropy, while the t-SNE representation suggests a higher value. This discrepancy underscores the need for profound research on appropriate GAN evaluation metrics for multivariate time series. It is however important to note that distances in t-SNE representations are not necessarily comparable due to nonlinear transformations. Using the conventional GAN training scheme, longer training could have resulted in convergence, however, harbors the risk of overfitting to the training data and thereby leading to phenomena like mode dropping. Artifacts such as small fluctuations in the time series unfortunately could not be diminished. Regularizing GANs is an ongoing research problem, and the behavior of GANs needs to be further understood.
@@
    In summary, our proposed Transformer-based Wasserstein GAN archictecture is a promising candidate for better representation learning of multivariate series and could have an impact to related domains such as data augmentation in various fields involving multivariate time series [29, 30, 14, 9, 6, 10, 34] or generating counterfactual explanations [20].
    Appendix 0.A Appendix
@@
    All Python scripts regarding the model architectures, training, loading the dataset, and evaluation are available at https://github.com/lscharwaechter/TransformerGAN.
    0.A.1 Methods for stabilizing GAN training
    Training GANs is notoriously hard. Common problems include mode collapse (”Helvetica scenario”), vanishing gradients, and furthermore there is no guarantee of convergence after a certain amount of training time. The following list gives an incomplete overview of different techniques to counteract these phenomena.
    Mode collapse
    • The generator must not trained too much without updating the discriminator, as this could result in mode collapse [12]
    • Apply batch normalization to all layers except the output layers of the generator and the discriminator to prevent the generator from mode collapse [31]
    • The Wasserstein loss supports the discriminator to reject samples the generator stabilizes on. So the generator has to learn something new [2]
    • Incorporating label information in the adversarial regularization helps to construct different modes [27]
    • Updating the generator not only using the current discriminator’s output but also the outputs of future discriminator versions (Unrolled GANs) inhibits the optimization for a single discriminator and forces the generator to diversify its output [28]
    • Modifying the discriminator to make decisions based on multiple samples of the same class (true/fake) at once helps to identify homogeneous packs as artificial [24]
    Vanishing gradients
    • Using a modified min-max loss, the generator is trained to maximize $log(D(G(z)))$ to prevent the GAN from vanishing gradients and getting stuck in the early stages of training [12]
    • The Wasserstein loss helps to prevent vanishing gradients by providing information even when the discriminator is trained to optimality [2]
    Convergence
    • Adding noise to the discriminator’s input helps to stabilize GAN training and improve convergence [1]
    • Regularizing the weight updates of the discriminator by using the norm of the gradients increases training stability [32]
    • Label noise, i.e. occasionally flipping the labels, balances the training and prevents the discriminator from being too strong and outperforming the generator [41]
    Practical considerations
    • Sampling from a Gaussian distribution instead of a Uniform distribution during training and afterwards interpolating via a circle from point A to point B instead of a straight line results in sharper samples [39]
    • Smoothing the labels when training the discriminator reduces the vulnerability to adversarial samples [33]
    • Normalized stochastic gradient descent ascent nSGDA is a simpler alternative to Adam and reaches similar or even better performance [17]
    Further regularization steps regarding GAN training can be found in [33].
    0.A.2 Additional Results
    Figure 3: Ten random examples of time series signals from the NATOPS validation set
    Figure 4: Errors of the Transformer model without GAN regularization (TAE)
    Figure 4: Errors of the Transformer model without GAN regularization (TAE)
    Figure 5: Errors of the convolutional model without GAN regularization (CAE)
    Figure 6: Ten different time series samples generated by TAE
    Figure 7: Ten different time series samples generated by CAE
    Figure 8: Error plots of the Transformer model regularized with the conventional GAN scheme (TAE-GAN)
    Figure 9: Ten different time series samples generated by TAE-GAN
    Figure 10: Error plots of the Transformer model regularized with the Wasserstein GAN scheme (TAE-WGAN)
    Figure 11: Ten different time series samples generated by TAE-WGAN
    Figure 12: Error plots of the convolutional model regularized with the convolutional GAN scheme (CAE-GAN)
    Figure 13: Ten different time series samples generated by CAE-GAN
    Figure 14: Error plots of the convolutional model regularized with the Wasserstein GAN scheme (CAE-WGAN)
    Figure 15: Ten different time series samples generated by CAE-WGAN

### v2

@@ 3.0.4 CAE:
    This model shows low similarity (DTW) and low diversity (Entropy) compared to the other models. The low diversity is further illustrated by the distribution in the t-SNE representation in Figure 3. Similar to TAE-GAN, the distribution lies exclusively within a new mode; there is no overlap with the modes of the dataset as the latent space was not regularized any further.
@@ 3.0.5 CAE-GAN:
    The Entropy score shows a high diversity compared to the other models, also confirmed by the distribution in the t-SNE plot (Figure 3). Here, the artificial time series lie within the different modes of the dataset, although not sufficiently distributed. The similarity score (DTW) is still low in an overall comparison, however slightly decreased compared to CAE.
@@ 3.0.6 CAE-WGAN:
    With an Entropy of $0.855$ this model shows the highest diversity, which is also confirmed by the distribution in the t-SNE plot (Figure 3). The generated time series are well balanced among the different modes of the dataset. Nevertheless, the similarity (DTW) remained low and even worsened compared to CAE and CAE-GAN.
@@ 4 Conclusion
    In this work, a Transformer-based model is proposed to generate multivariate time series signals within the support of the underlying distribution of an existing dataset. The Transformer architecture was augmented with a bottleneck to act as an autoencoder and was optimized using a conventional GAN and the Wasserstein GAN approach. Additionally, the model was compared to an existing approach that uses a convolutional autoencoder for local neighborhood generation of time series signals. The results show that the Transformer model was more suitable for stable GAN training. Using the Wasserstein GAN approach, the model outperformed the other models in terms of similarity to the dataset, however not all modes were sufficiently represented. The diversity is estimated to be rather low based on the multivariate Entropy, while the t-SNE representation suggests a higher value. This discrepancy underscores the need for profound research on appropriate GAN evaluation metrics for multivariate time series. It is however important to note that distances in t-SNE representations are not necessarily comparable due to nonlinear transformations. Using the conventional GAN training scheme, longer training could have resulted in convergence and higher similarity, however, harbors the risk of overfitting to the training data and thereby leading to phenomena like mode dropping. Artifacts such as small fluctuations in the time series unfortunately could not be diminished. Regularizing GANs is an ongoing research problem, and the behavior of GANs needs to be further understood.
@@
    In summary, our proposed Transformer-based Wasserstein GAN archictecture is a promising candidate for better representation learning of multivariate time series and could have an impact to related domains such as data augmentation in various fields involving multivariate time series [25, 26, 13, 8, 5, 9, 29] or generating counterfactual explanations [18].
    5 Code availability
@@
    All Python scripts regarding the model architectures, training, loading the dataset, and evaluation are available at
    https://github.com/lscharwaechter/TransformerGAN.